OpenStudy (jack1):

hi guys... bit of a harder Q is anyone has a sec? Could someone walk me through a laplace transform of : \[\Large f(t) = 5u(t) + 5e^{-5t}u(t) + sin(wt)u(t)\] or a laplace transform of anything? please and thank you.

OpenStudy (twopointinfinity):

We can help. Free Cupcakes. @DUDE..IM..A..DUCK @Mchilds15 @Dmb1215 @AngelWilliams16 @random231 @SGTC! @Ray_Rae1297 @Punkchick @CoolKatt16 @Kalob @brandonE

OpenStudy (anonymous):

I would help for a cupcake ;-; But i dont know how to do that yet ._. Sorry

OpenStudy (perl):

i think paul has some notes on laplace transformations

OpenStudy (mchilds15):

It's easy...

OpenStudy (unklerhaukus):

is \(u\) the unit heaviside step function ?

OpenStudy (jack1):

unsure @UnkleRhaukus ... it's literally just a q of many in my textbook... but thought they would indicate if it was heaviside

OpenStudy (unklerhaukus):

\[\Large f(t) = 5u(t) + 5e^{-5t}u(t) + \sin(\omega t)u(t)\\ \ \\\Large\mathcal L\{f(t)\}=\mathcal L\{5u(t) + 5e^{-5t}u(t) + \sin(\omega t)u(t)\}\\ \ \\ \Large F(s)=5\mathcal L\{u(t)\}+5\mathcal L\{e^{-5t}u(t)\}+\mathcal L\{\sin(\omega t)u(t)\} \]

OpenStudy (unklerhaukus):

so \(u(t)\) is just some arbitrary function?

OpenStudy (perl):

I dont think so It is called the unit step function http://math.bd.psu.edu/faculty/jprevite/251f11/250booklaplacept2.pdf

OpenStudy (jack1):

believe so, not told otherwise

OpenStudy (perl):

also called Heaviside function